# Loopy: Taming Audio-Driven Portrait Avatar with Long-Term Motion Dependency

### v2

@@ Abstract
With the introduction of diffusion-based video generation techniques, audio-conditioned human video generation has recently achieved significant breakthroughs in both the naturalness of motion and the synthesis of portrait details. Due to the limited control of audio signals in driving human motion, existing methods often add auxiliary spatial signals to stabilize movements, which may compromise the naturalness and freedom of motion. In this paper, we propose an end-to-end audio-only conditioned video diffusion model named Loopy. Specifically, we designed an inter- and intra-clip temporal module and an audio-to-latents module, enabling the model to leverage long-term motion information from the data to learn natural motion patterns and improving audio-portrait movement correlation. This method removes the need for manually specified spatial motion templates used in existing methods to constrain motion during inference. Extensive experiments show that Loopy outperforms recent audio-driven portrait diffusion models, delivering more lifelike and high-quality results across various scenarios.
@@ 1 Introduction
Due to the rapid advancements GAN and diffusion models in the field of video synthesis (Bar-Tal et al., 2024; Blattmann et al., 2023a; b; Guo et al., 2023; Zhou et al., 2022; Gupta et al., 2023; Wang et al., 2023; Ho et al., 2022; Brooks et al., 2022; Wang et al., 2020; Singer et al., 2022; Li et al., 2018; Villegas et al., 2022), human video synthesis (Siarohin et al., 2019; 2021; Xu et al., 2024b; Hu, 2024; Corona et al., 2024) has gradually approached the threshold of practical usability in terms of quality, attracting significant attention in recent years. Among these, zero-shot audio-driven portrait synthesis has seen an explosion of research (He et al., 2023; Tian et al., 2024; Xu et al., 2024a; Wang et al., 2024; Chen et al., 2024; Xu et al., 2024b; Stypulkowski et al., 2024) since around 2020, due to its ability to generate talking head videos with minimal barriers to entry. Starting last year, diffusion model techniques were introduced, with end-to-end audio-driven models (Tian et al., 2024; Xu et al., 2024a; Chen et al., 2024) demonstrating more vivid synthesis results compared to existing methods.
@@
However, due to the weak correlation between audio and portrait motion, end-to-end audio-driven methods typically introduce additional conditions related to spatial motion to ensure temporal stability in the synthesized videos. Conditions such as face locators and speed layers (Tian et al., 2024; Xu et al., 2024a; Chen et al., 2024) restrict the range and velocity of portrait movements, potentially reducing the expressiveness of the final output, as illustrated in Figure 1. While the introduction of preset motion templates may mitigate this issue, it also introduces complications related to template selection, audio-motion synchronization, and repetitive template movements. From a model perspective, these limitations also hinder the full potential of video diffusion models in generating vivid motion. This paper aims to address this issue by proposing an audio-only conditioned portrait diffusion model, enabling the model to learn natural motion patterns from data without the need for spatial templates.
@@
Figure 1: Visual comparisons with existing methods. Existing methods struggle to generate natural movements; compared to reference images, their motion, posture, and expressions often resemble the reference or remain nearly static due to the auxiliary spatial conditions. In contrast, Loopy effectively generates natural anthropomorphic movements solely from audio, including detailed head movements and facial expressions. Video results are provided in the supplementary materials.
We first attempted to remove structures such as the face locator and speed layer. This resulted in more frequent undesirable motion patterns in the synthesized videos, such as sudden temporal jitter, frame corruption, increased autoregressive degradation, and unnatural temporal details, leading to a decrease in overall quality. In practice, in current diffusion-based frameworks, the conditions influencing motion include not only audio (which maps to motion patterns in a many-to-many manner and is hard to fully define motions) but also motion frames. Motion frames provide appearance information from preceding clips, strongly influencing motion generation. However, current methods usually generate based on 4 motion frames from the last clip and over ten target frames from the current clip. At a typical 25 fps, this covers about 0.5 seconds in total, with motion frames covering only 0.2 seconds. Such short motion frames cause the model to primarily extract appearance information from the last clip rather than temporal motion information, such as motion style. For example, in the case of eye blinking, 0.2 seconds of preceding information is insufficient for the model to determine whether blinking should occur (it cannot know if blinking has occurred previously, making blinking a probabilistic event, which may lead to long periods without blinking in the generated video). When motion style is hard to determine by audio and motion frames, it exhibits randomness, necessitating additional guidance from spatial conditions such as face box and movement speed. We attempted to increase the length of motion frames and found that this approach can generate larger and more flexible movements to some extent, although it may lead to increased instability. This observation reminds us that increasing the temporal receptive field in an appropriate manner may help capture motion patterns more effectively and aid in generating natural motion. Additionally, current methods inject audio features into the diffusion model directly through cross-attention, which makes it difficult for the model to learn the relationship between audio and portrait motion, instead modeling the relationship between audio and all video pixels. This phenomenon is also mentioned in the Hallo (Xu et al., 2024a) work. The weak correlation between audio and portrait also increases the likelihood of generating random and unsatisfactory movements, making it difficult to generate natural motion that corresponds well with the audio.
@@
Based on the above observations and considerations, we propose Loopy, an end-to-end template-free audio-conditioned diffusion model for portrait video generation, leveraging long-term motion dependency to generate vivid portrait videos. Specifically: for temporal aspect: We designed inter- and intra-clip temporal modules. Motion frames are modeled with a separate temporal layer to capture inter-clip temporal relationships, while the original temporal module focuses on intra-clip temporal modeling. Additionally, we introduced a temporal segment module to the intra-clip layer, extending the receptive field to over 100 frames (covering approximately 5 seconds at 25 fps, 30 times the original). For audio aspect: We introduced the audio-to-motion latents module, which transforms audio and facial motion-related features (landmarks, head motion variance, expression motion variance) into motion latents based on a shared feature space. These latents are inserted into the denoising net as conditions. During testing, motion latents are generated using only audio. This approach allows weakly correlated audio to leverage strongly correlated motion conditions, enhancing the modeling of the relationship between audio and portrait motion. Extensive experiments validate that our design effectively improves the naturalness of motion and the robustness of video synthesis. In summary, our contributions include:
@@
(1) We propose a template-free audio-conditioned diffusion model for portrait video generation, featuring inter- and intra-clip temporal modules that leverage long-term motion dependency to learn natural motion patterns. Additionally, the audio-to-latents module enhances the correlation between audio and portrait motion by using strongly correlated conditions during training.
@@
(2) We validated the effectiveness of our method on public datasets and evaluated the model’s capabilities across various application scenarios (including multiple types of images and audio), demonstrating that our model achieves more vivid and stable synthesis results compared to existing methods.
@@ 2 Related Works
Audio-driven portrait video generation has attracted significant attention in recent years, with numerous works advancing the field. These methods can be categorized into GAN-based and diffusion-based approaches based on their video synthesis techniques.
GAN-based methods (Zhou et al., 2020; Prajwal et al., 2020; Zhang et al., 2023b; Liang et al., 2022) typically consist of two key components: an audio-to-motion model and a motion-to-video model. These models are usually implemented independently. For example, MakeItTalk (Zhou et al., 2020) uses an LSTM module to predict the landmark coordinates corresponding to the audio, and then a warp-based GAN model converts the landmark signals into video images. SadTalker (Zhang et al., 2023b) utilizes the existing FaceVid2Vid (Wang et al., 2021) method as the image synthesizer, employing ExpNet and PoseVAE to transform audio features into the inputs required by FaceVid2Vid, thereby completing the audio-to-video generation. With the introduction of diffusion techniques, some methods have implemented the audio-to-motion module using diffusion models while retaining the independent implementation of the motion-to-video module. For instance, GAIA (He et al., 2023) uses a VAE to represent motion as motion latents and implements a motion latents-to-video generation model. Additionally, it designs a diffusion model to achieve audio-to-motion latents generation, thereby enabling audio-to-video generation. DreamTalk (Ma et al., 2023), Dream-Talk (Zhang et al., 2023a), and VASA-1 (Xu et al., 2024b) propose similar ideas, using PIRender (Ren et al., 2021), FaceVid2Vid (Wang et al., 2021), and MegaPortrait (Drobyshev et al., 2022) as their motion-to-video models, respectively, and designing audio-to-motion representation models to complete the audio-to-portrait video generation process.
Apart from the above types, EMO Portrait (Tian et al., 2024) achieves audio-to-portrait video generation using a single diffusion model, replacing the two-stage independent design of the audio-to-motion module and the motion-to-video model. Hallo (Xu et al., 2024a), EchoMimic (Chen et al., 2024) and VExpress (Wang et al., 2024) have improved their audio-to-video modeling based on a similar audio-to-video diffusion framework. Although these end-to-end methods can generate vivid portrait videos, they need to introduce spatial condition module, like face locator and speed layer, to constrain head movements for stability, limiting the diversity of motion models in practical applications and hindering the full potential of diffusion models.
@@ 3 Method
Figure 2: The framework of Loopy. it removes the commonly used face locator and speed layer modules found in existing methods. Instead, it achieves flexible and natural motion generation through the proposed inter/intra-clip temporal layer and audio-to-latents module.
In this section, we will introduce our method, Loopy. First, we will provide an overview of the framework, including the input, output, and key components of Loopy. Second, we will focus on the design of the inter-/intra-temporal modules, including the temporal segment module. Third, we will detail the implementation of the audio condition module. Finally, we will describe the implementation details during the training and testing of Loopy.
@@ 3.1 Framework
Our method is built upon Stable Diffusion (SD) and uses its initialization weights. SD is a text-to-image diffusion model based on the Latent Diffusion Model (LDM) (Rombach et al., 2022). It employs a pretrained VQ-VAE (Kingma, 2013; Van Den Oord et al., 2017) $\mathcal{E}$ to transform images from pixel space to latent space. During training, images are first converted to latents, i.e., $z_{0}=\mathcal{E}(I)$. Gaussian noise $\epsilon$ is then added to the latents in the latent space based on the Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020) for $t$ steps, resulting in a noisy latent $z_{t}$. The denoising net takes $z_{t}$ as input to predict $\epsilon$. The training objective can be formulated as follows:
@@
where $\epsilon_{\theta}$ represents the denoising net, including condition-related attention modules, which is the main part Loopy aims to improve. $c$ represents the text condition embedding in SD, and in Loopy, it includes audio, motion frames, and other additional information influencing the final generation. During testing, the final image is obtained by sampling from Gaussian noise and removing the noise based on DDIM (Song et al., 2020) or DDPM.
@@
As demonstrated in Figure 2, in the Loopy, the inputs to the denoising net include noisy latents, i.e., the VQ-VAE encoded image latents $z_{t}$. Unlike original SD, where the input is a single image, here it is a sequence of images representing a video clip. The inputs also include reference latents $c_{ref}$ (encoded reference image latents via VQ-VAE), audio embedding $c_{audio}$ (audio features of the current clip), motion frames $c_{mf}$ (image latents of the M frames sequence from the preceding clips), and timestep $t$. During training, additional facial movement-related features are involved: $c_{lmk}$ (facial keypoint sequence of the current clip), $c_{mov}$ (head motion variance of the current clip), and $c_{exp}$ (expression variance of the current clip). The output is the predicted noise $\epsilon$. The denoising network employs a Dual U-Net architecture (Hu, 2024; Zhu et al., 2023). This architecture includes an additional reference net module, which replicates the original SD U-Net structure but utilizes reference latents $c_{ref}$ as input. The reference net operates concurrently with the denoising U-Net. During the spatial attention layer computation in the denoising U-Net, the key and value features from corresponding positions in the reference net are concatenated with the denoising U-Net’s features along the token dimension before proceeding with the attention module computation. This design enables the denoising U-Net to effectively incorporate reference image features from the reference net. Additionally, the reference net also takes motion frames latents $c_{mf}$ as input for feature extraction, allowing these features to be utilized in subsequent temporal attention computations.
3.2 Inter/Intra- Clip Temporal Layers Design
@@
Here, we introduce the design of the proposed inter-/intra-clip temporal modules. Unlike existing methods (Tian et al., 2024; Xu et al., 2024a; Chen et al., 2024; Wang et al., 2024) that process motion frame latents and noisy latents features simultaneously through a single temporal layer, Loopy employs two temporal attention layers: the inter-clip temporal layer and the intra-clip temporal layer. The inter-clip temporal layer first handles the cross-clip temporal relationships between motion frame latents and noisy latents, while the intra-clip temporal layer focuses on the temporal relationships within the noisy latents of the current clip.
@@
First, we introduce the inter-clip temporal layer, initially ignoring the temporal segment module in Figure 2. We first collect the $m$ image latents from the preceding clip, referred to as motion frames latents $c_{mf}$. Similar to $c_{ref}$, these latents are processed frame-by-frame through the reference network for feature extraction. Within each residual block, the features $x_{c_{mf}}$ obtained from the reference network are concatenated with the features $x$ from the denoising U-Net along the temporal dimension. To distinguish the types of latents, we add learnable temporal embeddings. Subsequently, self-attention is computed on the concatenated tokens along the temporal dimension, referred to as temporal attention. The intra-clip temporal layer differs in that its input does not include features from motion frames latents, it only processes features from the noisy latents of the current clip. By separating the dual temporal layers, the model can better handle the aggregation of different semantic temporal features in cross-clip temporal relationships.
@@
Due to the independent design of the inter-clip temporal layer, Loopy is better equipped to model the motion relationships among clips. To enhance this capability, we introduce the temporal segment module before $c_{mf}$ enters the reference network. This module not only extends the temporal range covered by the inter-clip temporal layer but also accounts for the variations in information due to the varying distances of different clips from the current clip, as illustrated in Figure 3. The temporal segment module divides the original motion frame into multiple segments and extracts representative motion frames from each segment to abstract the segment. Based on these abstracted motion frames, we recombine them to obtain motion frame latents for subsequent computations. For the segmentation process, we define two hyperparameters: stride $s$ and expand ratio $r$. The stride $s$ represents the number of abstract motion frames in each segment, while the expand ratio $r$ is used to calculate the length of the original motion frames in each segment. The number of frames in the $i$-th segment, ordered from closest to farthest from the current clip, is given by $s\times r^{(i-1)}$. For example, with a stride $s=4$ and an expand ratio $r=2$, the first segment would contain 4 frames, the second segment would contain 8 frames, and the third segment would contain 16 frames. For the abstraction process after segmentation, we default to uniform sampling within each segment. In the experimental section, we investigate different segmentation parameters and abstraction methods. Since segmentation and abstraction directly affect the learning of long-term motion dependency, different methods have a significant impact on the results. The output of the temporal segment module, $c^{o}_{mf}$, can be defined as:
$c^{o}_{mf,i}=c_{mf,\left\lfloor\frac{\text{start\_index}_{i}+\text{end\_index}_{i}}{2}\right\rfloor}$ (2)
[…]
The temporal segment module rapidly expands the temporal coverage of the motion frames input to the inter-clip temporal layer while maintaining acceptable computational complexity. For closer frames, a lower expansion rate retains more details, while for distant frames, a higher expansion rate covers a longer duration. This approach helps the model better capture motion style from long-term information and generate temporally natural motion without spatial templates.

### v3

@@ Abstract
With the introduction of the video diffusion model, audio-conditioned human video generation has recently achieved significant breakthroughs in both the naturalness of motion and the synthesis of portrait details. Due to the limited control of audio signals in driving human motion, existing methods often add auxiliary spatial signals, such as movement regions, to stabilize movements. However, this compromises the naturalness and freedom of motion. To address this issue, we propose an end-to-end audio-only conditioned video diffusion model named Loopy. Specifically, we designed two key modules: an inter- and intra-clip temporal module and an audio-to-latents module. These enable the model to better utilize long-term motion dependencies and establish a stronger audio-portrait movement correlation. Consequently, the model can generate more natural and stable portrait videos with subtle facial expressions, without the need for manually setting movement constraints. Extensive experiments show that Loopy outperforms recent audio-driven portrait diffusion models, delivering more lifelike and high-quality results across various scenarios. Video samples are available at this URL.
@@ 1 Introduction
Due to the rapid advancements of GAN and diffusion models in the field of video synthesis (Bar-Tal et al., 2024; Blattmann et al., 2023a; b; Guo et al., 2023; Zhou et al., 2022; Gupta et al., 2023; Wang et al., 2023; Ho et al., 2022; Brooks et al., 2022; Wang et al., 2020; Singer et al., 2022; Li et al., 2018; Villegas et al., 2022), human video synthesis (Siarohin et al., 2019; 2021; Xu et al., 2024b; Hu, 2024; Corona et al., 2024; Lin et al., 2024) has gradually approached the threshold of practical usability in terms of quality, attracting significant attention in recent years. Among these, zero-shot audio-driven portrait synthesis has seen an explosion of research (He et al., 2023; Tian et al., 2024; Xu et al., 2024a; Wang et al., 2024; Chen et al., 2024; Xu et al., 2024b; Stypulkowski et al., 2024) since around 2020, due to its ability to generate portrait videos with a low barrier to entry. Recently, diffusion model techniques have been introduced, with end-to-end audio-driven models (Tian et al., 2024; Xu et al., 2024a; Chen et al., 2024) demonstrating more vivid synthesis results compared to existing methods.
@@
However, due to the weak correlation between audio and portrait motion, end-to-end audio-driven methods typically introduce additional spatial conditions to restrict movement areas and ensure temporal stability in the synthesized videos. While the introduction of preset motion templates for spatial conditions may mitigate this issue, it also introduces several problems related to template selection, audio-template synchronization and repetitive movements. These spatial conditions, or auxiliary designs, such as face locators and speed layers (Tian et al., 2024; Xu et al., 2024a; Chen et al., 2024), restrict the range and velocity of portrait movements while also hindering the full potential of video diffusion models in generating vivid motion. This is because the model tends to follow given movement information during training rather than learning to generate natural movements from audio. As shown in Figure 1, the overall richness of portrait motion in existing methods is limited. This paper aims to address this issue by proposing an audio-only conditioned portrait diffusion model, enabling the model to learn natural motion patterns entirely from data without the need for spatial constraints.
@@
Figure 1: Visual comparisons with existing methods. Existing methods struggle to generate natural movements. Compared to reference images, their motion, posture, and expressions often resemble the reference or remain nearly static due to the auxiliary spatial conditions. In contrast, Loopy effectively generates natural movements solely from audio, including detailed head movements and facial expressions. Video samples are provided in the supplementary materials.
In the absence of spatial conditions, the input information is insufficient to accurately infer portrait movement trends, which leads to inconsistent and frequently changing motion patterns. We discovered that employing the CFG technique (Ho & Salimans, 2022) can stabilize videos by making them relatively static in comparison to the reference image. However, this approach is suboptimal because it fails to teach the model natural motion generation. It merely replicates expressions and poses from the reference image, ultimately resulting in a decline in overall quality. In diffusion-based frameworks, motion is influenced by both audio and frames from preceding clips, referred to as motion frames. Motion frames provide appearance information from preceding clips, which strongly influences motion generation. However, current methods typically use fewer than 5 motion frames, covering only about 0.2 seconds at 25 FPS. This brief duration causes the model to extract appearance information rather than temporal information, such as motion style. For instance, 0.2 seconds of preceding information is insufficient for the model to determine whether eye blinking should occur, making it a random event rather than a naturally generated expression. When motion style is difficult to determine from audio and motion frames, it exhibits randomness, necessitating additional guidance from spatial conditions such as face boxes and movement speed. We attempted to directly increase the length of motion frames and found that this can generate more dynamic facial movement details, although it is not stable. This suggests that appropriately increasing the temporal receptive field may help capture motion patterns more effectively and aid in generating natural motion.
Besides limited expressiveness caused by restricted movement, current methods also lack dynamic detail in facial movements, resulting in stiff and unnatural expressions beyond the mouth area, as shown in Figure 1. This issue may stem from the weak correlation between audio and portrait motion, which causes the model to focus more on modeling the relationship between audio and video pixels, including a significant amount of irrelevant background motion, rather than the relationship between audio and portrait motion. This phenomenon has also been noted in some studies (Xu et al., 2024a).
@@
Building on the above observations and considerations, we propose an end-to-end audio-conditioned diffusion model for portrait video generation, Loopy, leveraging long-term motion dependencies to generate vivid portrait videos. Specifically, in terms of the temporal aspect, we designed inter- and intra-clip temporal modules. Motion frames are modeled with a separate temporal layer to capture inter-clip relationships, whereas the original temporal module focuses on intra-clip modeling. Additionally, inter-clip temporal layers are equipped with a temporal segment module that extends the receptive field to over 100 frames (approximately 5 seconds at 25 fps, 30 times the original). These modules help the model better leverage long-term motion dependency information, allowing it to learn natural and stable motion patterns from data without the need for movement constraints. Furthermore, in terms of the audio aspect, we introduced the audio-to-latents module, which transforms audio and facial motion-related features (such as head movement variance and expression variance) into motion latents in a shared feature space. These latents are then inserted into the denoising network as conditions. During testing, motion latents are generated solely from audio. This approach allows weakly motion-correlated audio to leverage strongly correlated conditions, thereby enhancing the relationship between audio and portrait motion, and improving subtle facial expressions. Extensive experiments validate that our design effectively enhances both the naturalness of motion and the robustness of video synthesis across various types of input images and audio combinations. In summary, our contributions are as follows:
@@
(1) We propose Loopy, an audio-driven diffusion model for portrait video generation. It features two key components: inter- and intra-clip temporal modules designed to learn natural motion patterns from long-term dependencies, and an audio-to-latents module that enhances audio-portrait motion correlation by using strongly correlated conditions during training. Loopy is capable of generating vivid talking portrait videos with subtle facial details without relying on motion constraints or templates.
@@
(2) We validated the effectiveness of our method on public datasets and evaluated the model’s capabilities across various scenarios, including different types of input images and audio. The results demonstrate that Loopy achieves more lifelike and stable synthesis compared to existing methods.
@@ 2 Related Works
Audio-driven portrait video generation has attracted significant attention in recent years, with numerous works advancing the field. Most of these methods can be categorized into GAN-based and diffusion-based approaches according to their video synthesis techniques.
GAN-based methods (Zhou et al., 2020; Prajwal et al., 2020; Zhang et al., 2023b; Liang et al., 2022; Jiang et al., 2024) typically consist of two key components: an audio-to-motion model and a motion-to-video model. These models are usually implemented independently. For example, MakeItTalk (Zhou et al., 2020) uses an LSTM module to predict landmarks based on the input audio, and then a warp-based GAN model converts the landmarks into video. SadTalker (Zhang et al., 2023b) utilizes the existing FaceVid2Vid (Wang et al., 2021) method as the image synthesizer, employing ExpNet and PoseVAE to transform audio features into the inputs required by FaceVid2Vid, thereby completing the audio-to-video generation. With the introduction of diffusion techniques, some methods have implemented the audio-to-motion module using diffusion models while retaining the independent implementation of the motion-to-video module. For instance, GAIA (He et al., 2023) uses a VAE to represent motion as motion latents and implements a motion latents-to-video generation model. Furthermore, it designs a diffusion model to achieve audio-to-motion latents generation, thereby enabling audio-to-video generation. DreamTalk (Ma et al., 2023), Dream-Talk (Zhang et al., 2023a), and VASA-1 (Xu et al., 2024b) propose similar ideas, using PIRender (Ren et al., 2021), FaceVid2Vid (Wang et al., 2021), and MegaPortrait (Drobyshev et al., 2022) as their motion-to-video models, respectively, and designing audio-to-motion diffusion models to complete the audio-to-portrait video generation process. Additionally, some methods Ye et al. (2024) also incorporate NeRF for video rendering, combined with an audio-to-motion model to achieve audio-driven portrait video generation.
Apart from the above types, DiffusedHead (Stypulkowski et al., 2024) and EMO (Tian et al., 2024) achieves audio-to-portrait video generation using a single diffusion model, replacing the two-stage independent design of the audio-to-motion module and the motion-to-video model. Recent methods, such as Hallo (Xu et al., 2024a), EchoMimic (Chen et al., 2024) and VExpress (Wang et al., 2024), improve the audio-to-video modeling capabilities by introducing techniques like attention reweighting and spatial loss based on the end-to-end audio-to-video diffusion framework but still retain motion constraint conditions. Although these end-to-end methods can generate decent portrait videos, they need to introduce spatial condition module, like face locator and speed layer, to constrain portrait movements for stability, limiting the model’s ability to generate diverse motions in practical applications and hindering the full potential of diffusion models.
@@ 3 Method
Figure 2: The framework of Loopy. it removes the commonly used face locator and speed layer modules in existing methods. Instead, it achieves flexible and natural motion generation through the proposed inter/intra-clip temporal layers, temporal segment module (TSM) and audio-to-latents modules.
In this section, we first provide an overview of the Loopy framework, including its input, output and key designs. Second, we focus on the design of the inter/intra- temporal modules, including the temporal segment module. Third, we detail the implementation of the audio-to-latents module. Finally, we describe the implementation details for training and testing Loopy.
@@ 3.1 Framework
Our method is built upon Stable Diffusion (SD) and uses pretrained weights for initialization. SD is a text-to-image diffusion model based on the Latent Diffusion Model (LDM) (Rombach et al., 2022). It employs a pretrained VQ-VAE (Kingma, 2013; Van Den Oord et al., 2017) $\mathcal{E}$ to transform images from pixel space to latent space. During training, images are first converted to latents, i.e., $z_{0}=\mathcal{E}(I)$. Gaussian noise $\epsilon$ is then added to the latents in the latent space based on the Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020) for $t$ steps, resulting in a noisy latent $z_{t}$. The denoising net takes $z_{t}$ as input and predicts $\epsilon$. The training objective can be formulated as follows.
@@
where $\epsilon_{\theta}$ represents the denoising net, including condition-related modules, which is the main part Loopy aims to improve. $c$ represents the text condition embeddings in SD, and in Loopy, it is replaced by audio, motion frames, and other additional information influencing the final generation. During testing, the final image is obtained by sampling from gaussian noise and removing the noise based on DDIM (Song et al., 2020) or DDPM.
@@
As demonstrated in Figure 2, in the Loopy, the inputs to the denoising net include noisy latents $z_{t}$. Unlike the original SD, where the input is a single image, here the input is a sequence of images representing a video. The inputs also include reference latents $c_{ref}$ (encoded reference image via VQ-VAE), audio embeddings $c_{audio}$ (audio features of the current clip), motion frames $c_{mf}$ (image latents of the frames from the preceding clips) and timestep $t$. During training, additional facial movement-related features are involved, the head movement variance $c_{mov}$ and the expression variance $c_{exp}$ of the current clip. The output is the predicted noise $\epsilon$. The denoising network employs a dual U-Net architecture (Hu, 2024; Zhu et al., 2023). This architecture includes an additional reference network, which replicates the original SD U-Net structure but utilizes reference latents $c_{ref}$ as input. The reference network operates concurrently with the denoising U-Net. During the spatial attention layer computation in the denoising U-Net, the key and value features from corresponding positions in the reference network are concatenated with the denoising U-Net’s features along the spatial dimension before being processed by the attention module. This design enables the denoising U-Net to effectively incorporate reference image features from the reference network. Additionally, the reference network also takes motion frames latents $c_{mf}$ as input for feature extraction, allowing these features to be utilized in subsequent temporal attention computations.
3.2 Inter/Intra- Clip Temporal Module
@@
Here, we introduce the design of the proposed inter/intra- clip temporal modules. Unlike existing methods (Tian et al., 2024; Xu et al., 2024a; Chen et al., 2024; Wang et al., 2024) that process motion frame latents and noisy latent features simultaneously through a single temporal layer, Loopy employs two temporal attention layers, the inter-clip temporal layer and the intra-clip temporal layer. The inter-clip temporal layer first handles the cross-clip temporal relationships between motion frame latents and noisy latents, while the intra-clip temporal layer focuses on the temporal relationships within the noisy latents of the current clip.
@@
Inter/Intra- Temporal Lyaers. First, we introduce the inter-clip temporal layer, initially ignoring the temporal segment module in Figure 2. As shown in Figure 3, we first collect the image latents from the preceding clip, referred to as motion frames latents. Similar to $c_{ref}$, these latents are processed frame-by-frame through the reference network for feature extraction. Within each residual block, the motion frames latent features obtained from the reference network are concatenated with the noise latent features from the denoising U-Net along the temporal dimension. To distinguish the types of latents, we add learnable temporal embeddings. Subsequently, self-attention is computed on the concatenated tokens along the temporal dimension, i.e., temporal attention. The intra-clip temporal layer differs in that its input does not include features from motion frames latents, it only processes features from the noisy latents of the current clip. By using inter/intra-clip temporal layers, the model better handles the aggregation of semantic temporal features across clips.
@@
Temporal Segment Module. Due to the design of the inter-clip temporal layer, Loopy can better model the motion relationships among clips. To further enhance this capability, we introduce the temporal segment module before $c_{mf}$ enters the reference network. This module not only extends the temporal range covered by the inter-clip temporal layer but also extracts temporal information at different granularities based on the distance of each preceding clip from the current clip, as illustrated in Figure 3. The temporal segment module divides the original motion frame into multiple segments and extracts representative motion frames from each segment to abstract the segment. Based on these abstracted motion frames, we recombine them to obtain new motion frame latents, $c^{o}_{mf}$, for subsequent inter-clip temporal module computations. For the segmentation process, we define two hyperparameters, stride $s$ and expand ratio $r$. The stride $s$ represents the number of abstract motion frames in each segment, while the expand ratio $r$ is used to calculate the number of original motion frames covered in each segment. The number of frames in the $i$-th segment, ordered from closest to farthest from the current clip, is given by $s\times r^{(i-1)}$. For example, with a stride $s=4$ and an expand ratio $r=2$ (also our default setting, and the total segment number is set to 5), the first segment would contain 4 frames, the second segment would contain 8 frames, and the third segment would contain 16 frames. For the abstraction process after segmentation, we default to uniform sampling within each segment. In the experimental section, we investigate different segmentation parameters and abstraction methods. Different approaches significantly impact the results, as segmentation and abstraction directly affect the learning of long-term motion dependencies. The output of the temporal segment module, $c^{o}_{mf}$, can be defined as:
$c^{o}_{mf,i}$ $=c_{mf,\left[\sum_{j=0}^{k-1}r^{j}\cdot s+r^{k}\cdot(i\mod s)\right]}$ (2)
where $k=\left\lfloor\frac{i}{s}\right\rfloor$. $c^{o}_{mf,i}$ is the $i$-th element of the output and represents the $(i\mod s)$-th abstracted appearance information of the k-th segment.
The temporal segment module rapidly expands the temporal coverage of the motion frames input to the inter-clip temporal layer while maintaining acceptable computational complexity. For closer frames, a lower expansion rate retains more details, while for distant frames, a higher expansion rate covers a longer duration. This approach helps the model better capture motion style from long-term motion information and generate temporally natural motion without spatial constraints.